a spherical drop of radius 10 m has absorbed 40 electrons. What is the energy required to add 1 more electron?

Grade:12

1 Answers

Ashish Gupta IIT Roorkee
9 years ago
Energy = Potential * Charge = (Kq/r) * e = K (40e)(e)/ 10 = 4 Ke^{2})/ 10 = 4 Ke2